# Protecting Federated Learning from Extreme Model Poisoning Attacks via Multidimensional Time Series Anomaly Detection

### v2

@@ Abstract
1 INTRODUCTION
@@
This work introduces a novel pre-aggregation filter robust to untargeted model poisoning attacks. Notably, this filter $(i)$ operates without requiring prior knowledge or constraints on the number of malicious clients and $(ii)$ inherently integrates temporal dependencies. The FL server can employ this filter as a preprocessing step before applying any aggregation function, be it standard like FedAvg or robust like Krum or Bulyan. Specifically, we formulate the problem of identifying corrupted updates as a multidimensional (i.e., matrix-valued) time series anomaly detection task. The key idea is that legitimate local updates, resulting from well-calibrated iterative procedures like stochastic gradient descent (SGD) with an appropriate learning rate, show higher predictability compared to malicious updates. This hypothesis stems from the fact that the sequence of gradients (thus, model parameters) observed during legitimate training exhibit regular patterns, as validated in Section 4.2.
Inspired by the matrix autoregressive (MAR) framework for multidimensional time series forecasting [15], we propose the FLANDERS (Federated Learning meets ANomaly DEtection for a Robust and Secure) filter. The main advantage of FLANDERS over existing strategies like FLDetector [44] is its resilience to large-scale attacks, where $50\%$ or more FL participants are hostile. We attribute such a capability to two key factors: $(i)$ FLANDERS works without knowing a priori the ratio of corrupted clients, and $(ii)$ it embodies temporal dependencies between intra- and inter-client updates, quickly recognizing local model drifts caused by evil players. Below, we summarize our main contributions:
$\bullet$ We provide empirical evidence that the sequence of models sent by legitimate clients is more predictable than those of malicious participants performing untargeted model poisoning attacks.
$\bullet$ We introduce FLANDERS, the first pre-aggregation filter for FL robust to untargeted model poisoning based on multidimensional time series anomaly detection.
$\bullet$ We integrate FLANDERS into Flower, [^2] https://flower.dev/ a popular FL simulation framework for reproducibility.
$\bullet$ We show that FLANDERS improves the robustness of the existing aggregation methods under multiple settings: different datasets, client’s data distribution (non-iid), models, and attack scenarios.
$\bullet$ We publicly release all the implementation code of FLANDERS along with our experiments. [^3] https://anonymous.4open.science/r/flanders_exp-7EEB
The remainder of the paper is structured as follows. Section 2 covers background and preliminaries. In Section 3, we discuss related work. Section 4 and Section 5 describe the problem formulation and the method proposed. Section 6 gathers experimental results. Finally, we conclude in Section 7.
2 BACKGROUND AND PRELIMINARIES
2.1 Federated Learning
@@
(i) $S$ randomly selects a subset of clients ${\mathcal{C}}^{(t)}\subseteq\mathcal{C}$, so that $1\leq|{\mathcal{C}}^{(t)}|\leq K$, and sends them the current, global model $\bm{\theta}^{(t)}$. At $t=1$, $\bm{\theta}^{(1)}$ is randomly initalized. In the following, we assume that the number of clients picked at each round is constant and fixed, i.e., $|\mathcal{C}^{(t)}|=m,~{}\forall t\in\{1,2,\ldots,T\}$.
(ii) Each selected client $c\in{\mathcal{C}}^{(t)}$ trains its local model $\bm{\theta}_{c}^{(t)}$ on its own private data $\mathcal{D}_{c}$ by optimizing the following objective, starting from $\bm{\theta}^{(t)}$: $\bm{\theta}_{c}^{(t)}=\text{argmin}_{\bm{\theta}^{(t)}}\mathcal{L}_{c}(\bm{\theta}^{(t)};\mathcal{D}_{c})$. The value $\bm{\theta}_{c}^{(t)}$ is computed via gradient-based methods like stochastic gradient descent and sent to $S$.
(iii) $S$ computes $\bm{\theta}^{(t+1)}=\phi(\{\bm{\theta}_{c}^{(t)}~{}|~{}c\in\mathcal{C}^{(t)}\})$ as the updated global model, where $\phi:\mathbb{R}^{d^{m}}\mapsto\mathbb{R}^{d}$ is an aggregation function; for example, $\phi=\frac{1}{m}\sum_{c\in\mathcal{C}^{(t)}}\bm{\theta}_{c}^{(t)}$, i.e., FedAvg or alike [30].
2.2 The Attack Model: Federated Aggregation under Model Poisoning
The most straightforward aggregation function $\phi$ the server can implement is FedAvg, which computes the global model as the average of the local model weights received from clients. FedAvg is effective when all the FL participants behave honestly [18, 26, 31]. Instead, this work assumes that an attacker controls a fraction $b=\lceil r*K\rceil,r\in[0,1]$ of the $K$ clients, i.e., $0\leq b\leq K$, known as malicious.
@@ 2.2 The Attack Model: Federated Aggregation under Model Poisoning
Attacker’s Capability. Like Sybil attacks to distributed systems [20], the attacker can inject $b$ fake clients into the FL system or compromise $b$ honest clients. The attacker can arbitrarily manipulate the local models sent by malicious clients to the server $S$. More formally, let $x\in\mathcal{C}$ be one of the $b$ corrupted clients selected by the server on the generic $t$-th FL round; it first computes its legitimate local model $\bm{\theta}_{x}^{(t)}$ without modifying its private data $\mathcal{D}_{x}$; then it finds $\check{\bm{\theta}}_{x}^{(t)}$ by applying a post hoc perturbation $\bm{\varepsilon}\in\mathbb{R}^{d}$ to $\bm{\theta}_{x}^{(t)}$. For example, $\check{\bm{\theta}}_{x}^{(t)}=\bm{\theta}_{x}^{(t)}+\bm{\varepsilon}$, where $\bm{\varepsilon}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$ is a Gaussian noise vector. More advanced attack strategies have been designed, as discussed in Section 6.2.
Attacker’s Knowledge. We assume the attacker knows its controlled clients’ code, local training datasets, and local models. Moreover, we consider the worst-case scenario, where the attacker is omniscient, i.e., it has full knowledge of the parameters sent by honest parties. This allows the attacker to crafting malicious local models that are close to legitimate ones, thus increasing the probability of being chosen by the server for the aggregation.
Attacker’s Behavior. In each FL round, the attacker, similar to the server, randomly selects $b$ clients to corrupt out of the $K$ available. Any of these $b$ malicious clients that happen to be among those selected by the server will poison their local models using one of the strategies outlined in Section 6.2.
@@ 3 REALATED WORK
4 PROBLEM FORMULATION
4.1 Time Series of Local Models
@@
4.2 Predictability of Local Models Evolution: Legitimate vs. Malicious Clients
@@
To demonstrate this behavior, we consider two clients $i,j\in\mathcal{C}$. Client $i$ is assumed to be a legitimate participant, while client $j$ acts maliciously. Furthermore, both clients are selected by the server over a sequence of consecutive $T$ FL rounds to train a global model on the MNIST dataset. Therefore, we examine the local models sent to the server at each round $t\in\{1,2,\ldots,T\}$ by clients $i$ and $j$. Specifically, these are $d$-dimensional vectors of real-valued parameters $\bm{\theta}_{i}^{(t)},\bm{\theta}_{j}^{(t)}\in\mathbb{R}^{d}$, such that $\bm{\theta}_{i}^{(t)}=(\theta_{i,1}^{(t)},\ldots,\theta_{i,d}^{(t)})$ and $\bm{\theta}_{j}^{(t)}=(\theta_{j,1}^{(t)},\ldots,\theta_{j,d}^{(t)})$, respectively. Next, we calculate the average time-delayed mutual information (TDMI) for each pair of observed local models $(\bm{\theta}_{i}^{(t)},\bm{\theta}_{i}^{(t+\delta)})$ and $(\bm{\theta}_{j}^{(t)},\bm{\theta}_{j}^{(t+\delta)})$ across $T=50$ rounds, where $\delta>0$, for both client $i$ and $j$. This analysis aims to discern the time-dependent nonlinear correlation and the level of predictability between observations. We expect the TDMI calculated between pairs of models sent by the legitimate client $i$ to be higher than that computed between pairs sent by the malicious client $j$. Thus, the temporal sequence of local models observed from client $i$ is more predictable than that of client $j$.
Let $\bm{\theta}_{c}^{(t)}=(\theta_{c,1}^{(t)},\ldots,\theta_{c,d}^{(t)})$ denote the local model sent by the generic FL client $c\in\mathcal{C}$ as an instance of a $d$-dimensional vector-valued process. We define $\text{TDMI}(\bm{\theta}_{c}^{(t)},\bm{\theta}_{c}^{(t+\delta)})$ as follows:
@@
where $p(\cdot)$ is the probability density function. If each realization $\theta_{c,k}^{(t)}$ in the vector $\bm{\theta}_{c}^{(t)}$ is independent from each other, we can compute the average TDMI as follows [5]:
$\small\text{Avg}\Big{[}\text{TDMI}(\bm{\theta}_{c}^{(t)},\bm{\theta}_{c}^{(t+\delta)})\Big{]}=\frac{1}{d}\Bigg{[}\sum_{k=1}^{d}\text{TDMI}(\theta_{c,k}^{(t)},\theta_{c,k}^{(t+\delta)})\Bigg{]},$ (2)
where TDMI$(\theta_{c,k}^{(t)},\theta_{c,k}^{(t+\delta)})$ calculates TDMI for the univariate case. Indeed, the $d$ model parameters $\theta_{c,k}^{(\cdot)}$ should reasonably be independent. Under this assumption, the joint probability density function of the parameters factors into a product of individual probability density functions, forming a product measure on $d$-dimensional Euclidean space. Thus, according to Fubini’s theorem [12], the integral of each parameter will be independent of the others because the parameters are independent.
We consider the empirical distributions of the average TDMI between each pair of consecutive models $(\bm{\theta}_{c}^{(t)},\bm{\theta}_{c}^{(t+1)})$, sent by the legitimate client ($c=i$) and the malicious client ($c=j$) when running one of the four attacks described in Section 6.2 (see Fig. 1 in Appendix C). We call $\bar{\bm{\theta}_{i}}$ and $\bar{\bm{\theta}_{j}}$ the mean of these empirical distributions, respectively. Then, we run a one-tailed $t$-test against the null hypothesis $H_{0}:\bar{\bm{\theta}_{i}}=\bar{\bm{\theta}_{j}}$, where the alternative hypothesis is $H_{a}:\bar{\bm{\theta}_{i}}>\bar{\bm{\theta}_{j}}$. In all four cases, there is enough evidence to reject the null hypothesis at a confidence level $\alpha=0.01$ (see Tab. 4 in Appendix C).
This empirical evidence supports our claim and shows that the sequence of models sent by legitimate clients to the FL server is more predictable. Further discussion is provided in Appendix C.
4.3 Poisoned Local Models as Matrix-Based Time Series Outliers
@@
5 PROPOSED METHOD: FLANDERS
5.1 Matrix Autoregressive Model (MAR)
@@
where $||\cdot||_{\text{F}}$ indicates the Frobenius norm. Note that $l$ impacts only the size of the training set used to estimate the optimal coefficients $\hat{\bm{A}}$ and $\hat{\bm{B}}$; it does not affect the order of the autoregressive model, which will remain MAR($1$) and not MAR($l$). The optimal coefficients $\hat{\bm{A}}$ and $\hat{\bm{B}}$ can thus be estimated via alternating least squares (ALS) optimization [27]. Further details are provided in Appendix D.
5.2 MAR-based Anomaly Score
@@
MAR Estimation. At the first FL round ($t=1$), the server sends the initial global model $\bm{\theta}^{(1)}$ to the set of $m$ selected clients $\mathcal{C}^{(1)}$ and collects from them the $d\times m$ matrix of updated models $\bm{\Theta}_{1}$. Hence, it computes the new global model $\bm{\theta}^{(2)}=\phi(\{\bm{\theta}_{c}^{(1)}~{}|~{}c\in\mathcal{C}^{(1)}\})$, where $\phi=\text{Krum}$ or any other existing robust aggregation heuristic. For any other FL round $t>1$, the server can use the past $l>0$ historical observations $\bm{\Theta}_{t-l:t-1}$ to estimate the best MAR coefficients $\hat{\bm{\Omega}}=\{\hat{\bm{A}},\hat{\bm{B}}\}$ according to (5). In general, $11\leq l\leq t-$ 1; however, if we assume $l$ fixed at each round, the server will consider $\bm{\Theta}_{\max(1,t-l):t-1}$ past observations. Again, independently of the value of $l$, MAR will learn to predict the current matrix of weights only from the previously observed matrix. Therefore, each matrix used for training $\{\bm{\Theta}_{t-j-1}\}_{j=0}^{l-1}$ has the same size $d\times m$, as it contains exclusively the $m$ updates received at the round $t-j-1$. Of course, employing a higher order MAR($w$) model with $w>1$ would require extending each observed matrix to $d\times h$ ($m<h\leq K)$, as detailed in Section 4.1. This is to track the local updates sent by selected clients across multiple historical rounds.
@@
According to one of the filtering strategies discussed in Section 4.3, we retain only the $k$ clients with the smallest anomaly scores. The remaining $m-k$ clients are considered malicious; thus, they are discarded and do not contribute to the aggregation run by the server as if they were never selected.
@@
A comprehensive overview of how FLANDERS works along with its pseudocode is in Appendix E.
5.3 Computational Complexity Analysis

### v3

@@ Abstract
Index Terms:
Federated Learning, Anomaly Detection, Model Poisoning Attack
I Introduction
@@
This work introduces a novel pre-aggregation filter robust to untargeted model poisoning attacks. Notably, this filter $(i)$ operates without requiring prior knowledge or constraints on the number of malicious clients and $(ii)$ inherently integrates temporal dependencies. The FL server can employ this filter as a preprocessing step before applying any aggregation function, be it standard like FedAvg or robust like Krum or Bulyan. Specifically, we formulate the problem of identifying corrupted updates as a multidimensional (i.e., matrix-valued) time series anomaly detection task. The key idea is that legitimate local updates, resulting from well-calibrated iterative procedures like stochastic gradient descent (SGD) with an appropriate learning rate, show higher predictability compared to malicious updates. This hypothesis stems from the fact that the sequence of gradients (thus, model parameters) observed during legitimate training exhibit regular patterns, as validated in Section IV-B.
Inspired by the matrix autoregressive (MAR) framework for multidimensional time series forecasting [15], we propose the FLANDERS (Federated Learning meets ANomaly DEtection for a Robust and Secure) filter. The main advantages of FLANDERS over existing strategies like FLDetector [46] are its resilience to large-scale attacks, where $50\%$ or more FL participants are hostile, and the capability of working under realistic non-iid scenarios. We attribute such a capability to two key factors: $(i)$ FLANDERS works without knowing a priori the ratio of corrupted clients, and $(ii)$ it embodies temporal dependencies between intra- and inter-client updates, quickly recognizing local model drifts caused by evil players. Below, we summarize our main contributions:
  • (i)
    We provide empirical evidence that the sequence of models sent by legitimate clients is more predictable than those of malicious participants performing untargeted model poisoning attacks.
  • (ii)
    We introduce FLANDERS, the first pre-aggregation filter for FL robust to untargeted model poisoning based on multidimensional time series anomaly detection.
  • (iii)
    We integrate FLANDERS into Flower, [^2] https://flower.dev/ a popular FL simulation framework for reproducibility.
  • (iv)
    We show that FLANDERS improves the robustness of the existing aggregation methods under multiple settings: different datasets, client’s data distribution (non-iid), models, and attack scenarios.
  • (v)
    We publicly release all the implementation code of FLANDERS along with our experiments. [^3] https://anonymous.4open.science/r/flanders_exp-7EEB
The remainder of the paper is structured as follows. Section II covers background and preliminaries. In Section III, we discuss related work. Section IV and Section V describe the problem formulation and the method proposed. Section VI gathers experimental results. Finally, we conclude in Section VIII.
II Background and Preliminaries
II-A Federated Learning
@@
  1. (i)
    $S$ randomly selects a subset of clients ${\mathcal{C}}^{(t)}\subseteq\mathcal{C}$, so that $1\leq|{\mathcal{C}}^{(t)}|\leq K$, and sends them the current, global model $\bm{\theta}^{(t)}$. At $t=1$, $\bm{\theta}^{(1)}$ is randomly initalized.
  2. (ii)
    Each selected client $c\in{\mathcal{C}}^{(t)}$ trains its local model $\bm{\theta}_{c}^{(t)}$ on its own private data $\mathcal{D}_{c}$ by optimizing the following objective, starting from $\bm{\theta}^{(t)}$:
    $\bm{\theta}_{c}^{(t)}=\text{argmin}_{\bm{\theta}^{(t)}}\mathcal{L}_{c}(\bm{\theta}^{(t)};\mathcal{D}_{c}).$
    The value of $\bm{\theta}_{c}^{(t)}$ is computed via gradient-based methods (e.g., stochastic gradient descent) and sent to $S$.
  3. (iii)
    $S$ computes $\bm{\theta}^{(t+1)}=\phi(\{\bm{\theta}_{c}^{(t)}~{}|~{}c\in\mathcal{C}^{(t)}\})$ as the updated global model, where $\phi:\mathbb{R}^{d^{m}}\mapsto\mathbb{R}^{d}$ is an aggregation function; for example, $\phi=\frac{1}{m}\sum_{c\in\mathcal{C}^{(t)}}\bm{\theta}_{c}^{(t)}$, i.e., FedAvg or alike [31].
II-B The Attack Model: Federated Aggregation under Model Poisoning
The most straightforward aggregation function $\phi$ the server can implement is FedAvg, which computes the global model as the average of the local model weights received from clients. FedAvg is effective when all the FL participants behave honestly [18, 27, 32]. Instead, this work assumes that an attacker controls a fraction $b=\lceil r*K\rceil,r\in[0,1]$ of the $K$ clients, i.e., $0\leq b\leq K$, known as malicious. This is in contrast to previous works, where $r<0.5$. Below, we describe our attack model.
@@ II-B The Attack Model: Federated Aggregation under Model Poisoning
Attacker’s Capability. Like Sybil attacks to distributed systems [20], the attacker can inject $b$ fake clients into the FL system or compromise $b$ honest clients. The attacker can arbitrarily manipulate the local models sent by malicious clients to the server $S$. More formally, let $x\in\mathcal{C}$ be one of the $b$ corrupted clients selected by the server on the generic $t$-th FL round; it first computes its legitimate local model $\bm{\theta}_{x}^{(t)}$ without modifying its private data $\mathcal{D}_{x}$; then it finds $\check{\bm{\theta}}_{x}^{(t)}$ by applying a post hoc perturbation $\bm{\varepsilon}\in\mathbb{R}^{d}$ to $\bm{\theta}_{x}^{(t)}$. For example, $\check{\bm{\theta}}_{x}^{(t)}=\bm{\theta}_{x}^{(t)}+\bm{\varepsilon}$, where $\bm{\varepsilon}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$ is a Gaussian noise vector. More advanced attack strategies have been designed, as discussed in Section VI-B.
Attacker’s Knowledge. We assume the attacker knows its controlled clients’ code, local training datasets, and local models. Moreover, we consider the worst-case scenario, where the attacker is omniscient, i.e., it has full knowledge of the parameters sent by honest parties. This allows the attacker to choose the best parameters to inject into the FL protocol, e.g., by crafting malicious local models that are close to legitimate ones, thus increasing the probability of being chosen by the server for the aggregation.
Attacker’s Behavior. In each FL round, the attacker, similar to the server, randomly selects $b$ clients to corrupt out of the $K$ available. Any of these $b$ malicious clients that happen to be among those selected by the server will poison their local models using one of the strategies outlined in Section VI-B. Note that, unlike earlier studies, we do not assume that malicious clients are chosen in the initial round and remain constant throughout the training. We argue that this approach is more realistic, as legitimate clients can be compromised at any point in actual deployments. Therefore, an effective defense in FL should ideally exclude these clients as soon as they submit their first malicious model.
@@ III Related Work
IV Problem Formulation
IV-A Time Series of Local Models
@@
IV-B Predictability of Local Models Evolution: Legitimate vs. Malicious Clients
@@
To demonstrate this behavior, we consider two clients $i,j\in\mathcal{C}$. Client $i$ is assumed to be a legitimate participant, while client $j$ acts maliciously. Both are selected by the server over a sequence of consecutive $T$ FL rounds to train a global model on the MNIST dataset. Therefore, we examine the local models sent to the server at each round $t\in\{1,2,\ldots,T\}$ by clients $i$ and $j$. Specifically, these are $d$-dimensional vectors of real-valued parameters $\bm{\theta}_{i}^{(t)},\bm{\theta}_{j}^{(t)}\in\mathbb{R}^{d}$, such that $\bm{\theta}_{i}^{(t)}=(\theta_{i,1}^{(t)},\ldots,\theta_{i,d}^{(t)})$ and $\bm{\theta}_{j}^{(t)}=(\theta_{j,1}^{(t)},\ldots,\theta_{j,d}^{(t)})$, respectively. Next, we calculate the average time-delayed mutual information (TDMI) for each pair of observed local models $(\bm{\theta}_{i}^{(t)},\bm{\theta}_{i}^{(t^{\prime})})$ and $(\bm{\theta}_{j}^{(t)},\bm{\theta}_{j}^{(t^{\prime})})$ across $T=50$ rounds, where $\delta>0$, for both client $i$ and $j$. This analysis aims to discern the time-dependent nonlinear correlation and the level of predictability between observations. We expect the TDMI calculated between pairs of models sent by the legitimate client $i$ to be higher than that computed between pairs sent by the malicious client $j$. Thus, the temporal sequence of local models observed from client $i$ is more predictable than that of client $j$.
Let $\bm{\theta}_{c}^{(t)}=(\theta_{c,1}^{(t)},\ldots,\theta_{c,d}^{(t)})$ denote the local model sent by the generic FL client $c\in\mathcal{C}$ as an instance of a $d$-dimensional vector-valued process. We define $\text{TDMI}(\bm{\theta}_{c}^{(t)},\bm{\theta}_{c}^{(t^{\prime})})$ as follows:
@@
where $p(\cdot)$ is the probability density function and $t^{\prime}=t+\delta$. If each realization $\theta_{c,k}^{(t)}$ in the vector $\bm{\theta}_{c}^{(t)}$ is independent from each other, we can compute the average TDMI as follows [5]:
$\small\text{Avg}\Big{[}\text{TDMI}(\bm{\theta}_{c}^{(t)},\bm{\theta}_{c}^{(t^{\prime})})\Big{]}=\frac{1}{d}\Bigg{[}\sum_{k=1}^{d}\text{TDMI}(\theta_{c,k}^{(t)},\theta_{c,k}^{(t^{\prime})})\Bigg{]},$ (3)
where TDMI$(\theta_{c,k}^{(t)},\theta_{c,k}^{(t^{\prime})})$ calculates TDMI for the univariate case. Indeed, the $d$ model parameters $\theta_{c,k}^{(\cdot)}$ should reasonably be independent. Under this assumption, the joint probability density function of the parameters factors into a product of individual probability density functions, forming a product measure on $d$-dimensional Euclidean space. Thus, according to Fubini’s theorem [12], the integral of each parameter will be independent of the others because the parameters are independent.
To mimic the behavior of a hypothetical optimal FLANDERS filter, and therefore avoid the propagation of poisoned local models sent by client $j$ across the $T$ rounds, we assume that, at each round $t$, the global model from which both clients start their local training process is polished from any malicious updates received at the previous round $t-1$. Next, we calculate the average time-delayed mutual information (TDMI) for each pair of observed local models $(\bm{\theta}_{i}^{(t)},\bm{\theta}_{i}^{(t^{\prime})})$ and $(\bm{\theta}_{j}^{(t)},\bm{\theta}_{j}^{(t^{\prime})})$ across $T=50$ rounds, where $t^{\prime}>t$, for both client $i$ and $j$. Firstly, we consider the special case where $t^{\prime}=t+1$ and compute the average TDMI between each pair of consecutive local models sent by the legitimate client and the malicious client, when this runs one of the four attacks considered in this work, namely GAUSS, LIE, OPT, and AGR-MM presented in Section VI-B. In Figure 1, we plot the empirical distributions of the observed average TDMI for the legitimate and malicious clients.
[…]
TABLE I: One-tailed $t$-test against the null hypothesis $H_{0}:\bar{\bm{\theta}_{i}}=\bar{\bm{\theta}_{j}}^{atk}$. Each cell contains the $p$-value for the statistical test corresponding to a specific attack. In all four cases, there is enough evidence to reject the null hypothesis at a confidence level $\alpha=0.01$ ($p$-value $\ll 0.01$).
$H_{0}:\bar{\bm{\theta}_{i}}=\bar{\bm{\theta}_{j}}^{atk}$
GAUSS $5.67*10^{-38}$
LIE $8.33*10^{-5}$
OPT $6.70*10^{-18}$
AGR-MM $4.20*10^{-12}$
IV-C Poisoned Local Models as Matrix-Based Time Series Outliers
@@
V Proposed Method: FLANDERS
V-A Matrix Autoregressive Model (MAR)
@@
where $||\cdot||_{\text{F}}$ indicates the Frobenius norm. Note that $l$ impacts only the size of the training set used to estimate the optimal coefficients $\hat{\bm{A}}$ and $\hat{\bm{B}}$; it does not affect the order of the autoregressive model, which will remain MAR($1$) and not MAR($l$). The optimal coefficients $\hat{\bm{A}}$ and $\hat{\bm{B}}$ can thus be estimated via alternating least squares (ALS) optimization [28]. Further details are provided in Appendix XI.
V-B MAR-based Anomaly Score
@@
MAR Estimation. At the first FL round ($t=1$), the server sends the initial global model $\bm{\theta}^{(1)}$ to the set of $m$ selected clients $\mathcal{C}^{(1)}$ and collects from them the $d\times m$ matrix of updated models $\bm{\Theta}_{1}$. Hence, it computes the new global model $\bm{\theta}^{(2)}=\phi(\{\bm{\theta}_{c}^{(1)}~{}|~{}c\in\mathcal{C}^{(1)}\})$, where $\phi=\text{Krum}$ or any other existing robust aggregation heuristic. For any other FL round $t>1$, the server can use the past $l>0$ historical observations $\bm{\Theta}_{t-l:t-1}$ to estimate the best MAR coefficients $\hat{\bm{\Omega}}=\{\hat{\bm{A}},\hat{\bm{B}}\}$ according to (6). In general, $11\leq l\leq t-$ 1; however, if we assume $l$ fixed at each round, the server will consider $\bm{\Theta}_{\max(1,t-l):t-1}$ past observations. Again, independently of the value of $l$, MAR will learn to predict the current matrix of weights only from the previously observed matrix. Therefore, each matrix used for training $\{\bm{\Theta}_{t-j-1}\}_{j=0}^{l-1}$ has the same size $d\times m$, as it contains exclusively the $m$ updates received at the round $t-j-1$. Of course, employing a higher order MAR($w$) model with $w>1$ would require extending each observed matrix to $d\times h$ ($m<h\leq K)$, as detailed in Section IV-A. This is to track the local updates sent by selected clients across multiple historical rounds.
@@
According to one of the filtering strategies discussed in Section IV-C, we retain only the $k$ clients with the smallest anomaly scores. The remaining $m-k$ clients are considered malicious; thus, they are discarded and do not contribute to the aggregation run by the server as if they were never selected.
@@
V-C A Step-by-Step Example
In Figure 2, we depict how FLANDERS computes the anomaly score vector $\bm{s}^{(t)}$ at the generic FL round $t$. Specifically, the server $S$ uses its current MAR($1$) forecasting model $f$ whose best parameters $\hat{\bm{\Omega}}=\{\hat{\bm{A}},\hat{\bm{B}}\}$ are estimated from $l$ previous historical observations of local models received in the previous rounds $t-l,\ldots,t-1$. It applies $f$ to the previously observed matrix $\bm{\bm{\Theta}}_{t-1}$ to get the next predicted matrix of local models $\bm{\hat{\bm{\Theta}}}_{t}$, i.e., $\bm{\hat{\bm{\Theta}}}_{t}=f(\bm{\bm{\Theta}}_{t-1};\hat{\bm{\Omega}})=\hat{\bm{A}}\bm{\bm{\Theta}}_{t-1}\hat{\bm{B}}$. Then, it compares this predicted matrix $\bm{\hat{\bm{\Theta}}}_{t}$ with the actual matrix of local updates received $\bm{\bm{\Theta}}_{t}$. The final anomaly score vector is calculated by measuring the distance $\delta$ between each column of those two matrices, according to Eq. 4.
It is worth remarking that, in general, only some clients are selected at every round. In particular, if a client $c_{\text{new}}$ – whether it is honest or malicious – is selected by the server for the first time at round $t$, the MAR forecasting model $f$ will not be able to make any prediction for it due to a cold start problem (i.e., the predicted matrix $\bm{\hat{\bm{\Theta}}}_{t}$ will not contain a column corresponding to $c_{\text{new}}$). In such a case, we must adopt a fallback strategy. Without any historical information for a client, the most sensible thing to do is to compute the distance $\delta$ between the local model update it sent and the current global model.
Figure 2: Overview of FLANDERS.
[…]
It is worth noting that $\mathcal{C}^{(1)}\cap\mathcal{C}^{(2)}=\{c_{3},c_{9}\}$; the other two clients, $c_{1}$ and $c_{6}$, are considered “cold-start” since they are selected for the first time or, in any case, they do not appear in the historical observations used by MAR for predictions (i.e., in the previous $w=1$ matrices). Thus, we calculate the following anomaly scores, according to Eq. 5:
To overcome this problem, FLANDERS replaces the local model marked as suspicious $\bm{\theta}_{3}^{(2)}$ with either one of the previously observed models from the same client that is supposedly legitimate or the current global model. Since, in this example, we assume $c_{3}$ has been malicious from the first round, we use the latter approach. Specifically, we change $\bm{\Theta}_{2}$ with $\bm{\Theta}^{\prime}_{2}=[\bm{\theta}_{1}^{(2)},\underline{\bm{\theta}^{(2)}},\bm{\theta}_{6}^{(2)},\bm{\theta}_{9}^{(2)}]$, where $\bm{\theta}^{(2)}$ substitutes $\bm{\theta}_{3}^{(2)}$. As discussed at the end of Section V-B, this fix allows FLANDERS to work even when a malicious client is picked for two or more rounds consecutively.
Therefore, anomaly scores are updated as follows:
V-D FLANDERS’ Pseudocode
Algorithm 1 FLANDERS-Server
[…]
2:$\mathcal{C}^{(t)}\leftarrow$ sample a subset of $m$ clients from $\mathcal{C}$
3:Send the global model $\bm{\theta}^{(t)}$ to every $c\in\mathcal{C}^{(t)}$
[…]
6:$\mathcal{C}^{(t)}_{*}\leftarrow\texttt{{FLANDERS}}(\bm{\Theta}_{t-l:t-1},\mathcal{C}^{(t)},k,w,N)$EndIf
[…]
Algorithm 2 FLANDERS-Filter
[…]
1:$\hat{\bm{\Theta}}_{t}\leftarrow\texttt{MAR}(\bm{\Theta}_{t-l:t-1},w,N)$
2:$\bm{s}^{(t)}\leftarrow\delta(\bm{\Theta}_{t},\hat{\bm{\Theta}}_{t})$triangleright Anomaly score vector
3:$\mathcal{C}^{(t)}_{*}\leftarrow\{c\in\mathcal{C}^{(t)}~{}|~{}s_{c}^{(t)}\leq s_{k}^{(t)}\}$triangleright $s^{(t)}_{k}$ is the $k$-th smallest anomaly score
4:return $\mathcal{C}^{(t)}_{*}$EndProcedure
[…]
V-E Computational Complexity Analysis